OpenStudy (anonymous):

If I was to create my own function, and then assign any number to x. How would I explain whether f(h(x)) and h(f(x)) will always result in the same number?

OpenStudy (anonymous):

@ParthKohli could you help?

Parth (parthkohli):

Hi, have you heard of an inverse function?

OpenStudy (anonymous):

would it help if I went ahead and made a function?

OpenStudy (anonymous):

i have.

Parth (parthkohli):

Yes. Two inverse functions are just like that.

OpenStudy (anonymous):

you're saying f(h(x)) and h(f(x)) are the inverse of each other?

OpenStudy (anonymous):

Let's say my function was this----> F(x)=7x-2

OpenStudy (anonymous):

you'll get your medals, I appreciate the help.

Parth (parthkohli):

Yes, that's how I mean it. For example, multiplying and dividing by the same number.

Parth (parthkohli):

Eh, no medals needed... x/2 and 2x are inverse functions.

OpenStudy (anonymous):

wow, you the man.

Parth (parthkohli):

?

OpenStudy (anonymous):

okay I see what you mean.

OpenStudy (anonymous):

I'd need to create an inverse function? In order to show this.

Parth (parthkohli):

Yes, any two functions that are inverses of each other would work.

OpenStudy (anonymous):

how would that show whether f(h(x)) and h(f(x)) will always result in the same number? @ParthKohli

Parth (parthkohli):

OK, let's show it. Let's say that h(x) = x * 1/2 and f(x) = 2 * x What is f(h(x))?

OpenStudy (anonymous):

I'd have to solve for X, solve each one individually, and then do the same for h(f(x)) and Voila, that's my answer?

OpenStudy (anonymous):

@ParthKohli

terenzreignz (terenzreignz):

Hello there, mate~ Mind if I watch you tackle this minor fiasco? ^_^

Parth (parthkohli):

Yes, that is right.

OpenStudy (anonymous):

My fiasco, is your fiasco...we've just become brothers @terenzreignz

Parth (parthkohli):

Can you calculate f(h(x))?

OpenStudy (anonymous):

@ParthKohli you rock, thanks. :)

OpenStudy (anonymous):

I don't know how to calculate F(h(x)) :/

Parth (parthkohli):

Eh, not necessarily... but thanks! You know what f(h(x)) means, right?

terenzreignz (terenzreignz):

Brothers, are we? Well then, for starters, it's TJ :) Perhaps we might need a little review of what functions are and their not-so complex complexities?

Parth (parthkohli):

We're doing the thing "h" to f(x). h(x) = 1/2 * x Doing the thing "h" means that you take something and halve it.

Parth (parthkohli):

So, you are going to take f(x) and halve it. f(x) = 2 * x right? What is the half of 2 * x?

Parth (parthkohli):

In \(h(\color{blue}{f(x}))\), that is.

OpenStudy (anonymous):

The complexity of their level of complex reaches into layers and layers of complexity until you find yourself, much like the panda that learned Kung-Fu, in a void with no escape...lol I'll take all the help I can get @terenzreignz

OpenStudy (anonymous):

would it be x/2? O.o @ParthKohli

Parth (parthkohli):

\[h(f(x))\]\[= h(\color{blue}{2\times x})\]\[= \dfrac{1}{2}\times \color{blue}{2 \times x}\]

OpenStudy (anonymous):

arhhhg, sorry man...Math is dead to me. :/ I suck at it

OpenStudy (anonymous):

that's the final product? O.o

Parth (parthkohli):

Nope! What is 1/2 * 2 * x?\[\left(\dfrac{1}{2}\times 2\right) \times x\]

OpenStudy (anonymous):

you see to me that looks like a bunch of mumbo-jumbo...I would say 1 * x

Parth (parthkohli):

Yes! That is right. 1 * x is just x, right?

OpenStudy (anonymous):

because .5, or half of 1 multiplied by 2, is...1

OpenStudy (anonymous):

yah, it would just be x.

Parth (parthkohli):

Yup, you got it. How about f(h(x))? Since the question is looking for both h(f(x)) and f(h(x))...

OpenStudy (anonymous):

ok so the final product for hfx would be just X?

OpenStudy (anonymous):

and I would say...fhx would be 2x

Parth (parthkohli):

The final product of h(f(x)) is just x, yes. How can f(h(x)) be 2x...? Ideally, they should both be x as they result in the same number. :P

terenzreignz (terenzreignz):

May I give it a shot? Watching is boring. :3

OpenStudy (anonymous):

lol it's all you man

OpenStudy (anonymous):

are you solving for F(x)=7x-2? that function :$

terenzreignz (terenzreignz):

All right, Jack.. first, no tagging me all the time, I'm already on this question, I got a name, it's TJ (not really, LOL), so use it. Now, to more important matters... a function is like a process... it takes an input, usually called x, and sort of... *does something* to it. It can be anything really, from doubling... \[\Large f(\color{red}x) = 2\color{red}x\] To squaring... \[\Large f(\color{red}x)= \color{red}x^2\] Or even just raising by 4... \[\Large f(\color{red}x) = \color{red}x+4\] Catch me so far?

OpenStudy (anonymous):

so far TJ, I follow.

terenzreignz (terenzreignz):

Good. Now, x is technically just a placeholder... see that first function, the 'doubling' function? We can replace x with virtually anything (no wisecrack comments please :3 ) And the function will have a value...or output. Say, \[\Large f(\color{red}2) = 2(\color{red}2)=4\] Makes sense, yeah? Doubling 2 gives us 4. Still catching me so far?

OpenStudy (anonymous):

oooooh

OpenStudy (anonymous):

it clicked... -.-

OpenStudy (anonymous):

that's the point of the question!

OpenStudy (anonymous):

that if the answer always ends up the same, which it will, then you can literally put anything in X.

OpenStudy (anonymous):

It makes sense now..

terenzreignz (terenzreignz):

Focus. Try not to jump to conclusions. Arriving at the correct conclusion through the wrong logic occurs quite often in Maths, so stay on track. Now, when I said we can replace x with *anything* I mean even other functions. Suppose we have two functions. The doubling function\[\Large f(\color{red}x)=2\color{red}x\] and the "raise-by-four" function:\[\Large g(\color{red}x)= \color{red}x+4\] Still with me?

OpenStudy (anonymous):

yah.

terenzreignz (terenzreignz):

And just to illustrate, with the doubling function, the answer will NOT always be the same regardless of what we replace x with. Try doubling 2 and 3. You get 4 and 6, which aren't the same. So.. now...

terenzreignz (terenzreignz):

Enter a term. It's called "Function composition" It's when you apply a function to x, and then apply another function, NOT to x, but to the result of the first function. ^That may have been a lot to take in, so read through it, devour it, digest it. And then I'll illustrate an example.

OpenStudy (anonymous):

Digested *yum*

terenzreignz (terenzreignz):

Right. So what if I ask you for a function that FIRST doubles a number, and then adds four to it? f(x) was our doubling function, and g(x) was our "raise-by-four" function. The composition function is set up as follows: First you double:\[\Large f(\color{red}x)\]... and then you raise it by four: \[\Large g\left(\color{red}{f(x)}\right)\] ^ A bit tricky to get, so again, read thoroughly.

terenzreignz (terenzreignz):

Ready? What is this going to look like, I wonder...

OpenStudy (anonymous):

I'm honestly a bit confused at this point..

terenzreignz (terenzreignz):

Okay, it's good that you tell me in advance. If I ask you to double 5 and then add 4 to it, what would be the answer? ^_^

OpenStudy (anonymous):

14

terenzreignz (terenzreignz):

Great. So you can actually (albeit not consciously, apparently) compose functions. Now to the more technical stuffs... f(x) is the doubling function, so we apply it first, right?

OpenStudy (anonymous):

I would say so.

terenzreignz (terenzreignz):

Right, and whatever the output (represented by f(x) itself) we apply g(x) TO it. Still following? ^_^

OpenStudy (anonymous):

so far lol I follow

OpenStudy (anonymous):

you just left </3...

terenzreignz (terenzreignz):

When you do return... So, basically, we apply the function g to f(x). Like so: \[\Large g(\color{red}{f(x)})\] Like a mutant function of sorts. Now, I'd just like to recall that \[\Large g(\color{red}x)=\color{red}x+4\] Notice that whatever is in the parentheses beside g, it simply takes the place of all the x's in the function. Like so: \[\Large g(\color{red}y)=\color{red}y+4\]\[\Large g(\color{red}{a+b})=\color{red}{a+b}+4 \] And so on... still catching me?

OpenStudy (anonymous):

still catching

OpenStudy (anonymous):

I really have to solve this problem soon..

terenzreignz (terenzreignz):

So, what about \[\Large g(\color{red}{f(x)}) \] ? By that logic, you'd probably think that it should equal\[\Large g(\color{red}{f(x)})= \color{red}{f(x)} +4\] right?

OpenStudy (anonymous):

I'd say so :)

terenzreignz (terenzreignz):

And you'd be right. And remember that f(x) is simply 2x. So replace again, accordingly, and you get \[\Large g(f(\color{red}x))=2\color{red}x + 4 \]

OpenStudy (anonymous):

so g=4 and f(x)=2x

OpenStudy (anonymous):

hence, why they're equal :)

OpenStudy (anonymous):

makes more and more sense

terenzreignz (terenzreignz):

No.... g(x) = x+4 f(x) = 2x When you have g(f(x)), simply replace the x in g(x) with f(x), like so: g(f(x)) = f(x) + 4 And remember that f(x) was originally 2x, so g(f(x)) = 2x + 4

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so with that in mind...could you help me solve the question? O.o

OpenStudy (anonymous):

I have to create my own function, assign any number to x, explain whether f(h(x)) and h(f(x)) will always result in the same number. ...which they will

OpenStudy (anonymous):

I just need help explaining how lol

terenzreignz (terenzreignz):

Yes, well, there are certain pairs of functions that have a unique property... that is, they cancel each other out. These functions are called inverses. Think about it... it's like an 'undo' Like when doubling, to undo it, you halve. When squaring, to undo it, you get a square root. When adding four, to undo it, simply take away four. Get the picture?

terenzreignz (terenzreignz):

OH, is THAT the question? I thought you were asked to find a pair of functions where f(h(x)) is always equal to h(f(x)) -_-

OpenStudy (anonymous):

lol nope, that's the question I posed in the beggining.

OpenStudy (anonymous):

you have helped though, in my understanding of functions.

terenzreignz (terenzreignz):

Well then, that simplifies things somewhat. You still have to learn about compositions, though. Having said that, have you made up your own function?

terenzreignz (terenzreignz):

Are you already given a function?

OpenStudy (anonymous):

I'll make own now... f(x)=4x-1

OpenStudy (anonymous):

is that okay for a function?

terenzreignz (terenzreignz):

Of course. Although, from the looks of it, you have to make up another one. Unless you were already given another function... h(x)?

OpenStudy (anonymous):

sooo... h(x)=4x-1?

OpenStudy (anonymous):

it says to create your own function..

terenzreignz (terenzreignz):

It does... but does it say just one? Perhaps you need to make up two functions. sure they can be the same function as you have just done now, but that's boring, think of something else. (NOTE, this is IF you weren't given a specific function for h or f)

OpenStudy (anonymous):

just one